OpenStudy (anonymous):

WHICH OF THE FOLLOWING IS STANDARD DEVIATION OF THE SAMPLE SHOWN HERE ?? 17,18,19,20,21

OpenStudy (mathstudent55):

Are there choices?

OpenStudy (anonymous):

NO THE PROBLEM

OpenStudy (mathstudent55):

The numbers given are the sample?

OpenStudy (anonymous):

YES

OpenStudy (mathstudent55):

The question asks "which of the following". That means there must be a list of choices to choose from. Where are the choices?

OpenStudy (amistre64):

even without the options, we can still calculate the sd

OpenStudy (anonymous):

OH OK A.19 B.2 C.2.5 D.^2 E.^25

OpenStudy (amistre64):

we need to know the mean of the sample; then we can find the squared differences in order to determine the sample deviation from.

OpenStudy (mathstudent55):

Do you know how to calculate the standard deviation of a sample?

OpenStudy (anonymous):

no

OpenStudy (mathstudent55):

First, calculate the mean of the sample. This is just the arithmetic average. Add all numbers and divide by the number of numbers.

OpenStudy (kirbykirby):

yup and then the sample standard deviation is: \[\large \sqrt{\frac{1}{n-1}\sum_{i=1}^n(x_i-\bar{x})^2} \] where \(\bar{x}\) is the mean you find, and \(n\) is the number of data points in your sample (in your case, \(n=5\))

OpenStudy (kirbykirby):

If the notation is confusing, this is just telling you to find \[\sqrt{\frac{1}{5-1}[(17-\bar{x})^2+(18-\bar{x})^2+(19-\bar{x})^2+(20-\bar{x})^2+(21-\bar{x})^2}]\]